OpenStudy (anonymous):

The speed that a car is going on a dry road when the driver puts on the breaks can be estimated by the equation s=√24d , where d is the distance that the car skids in feet. A car that was in an accident skidded 214 ft. Find the speed, in miles per hour, that a car was going before the accident occurred. Round your answer to the nearest tenth.

OpenStudy (anonymous):

I understand this is not your question http://openstudy.com/study#/updates/4ff133a6e4b03c0c488a25e3 but you can go to it and plug your problem into it

OpenStudy (anonymous):

okay i will try

OpenStudy (phi):

they give you an equation \[ s = \sqrt{24d} \] d is the distance that the car skids A car skidded 214 ft. what number should you use for d ?

OpenStudy (anonymous):

24 @phi

OpenStudy (phi):

A car skidded 214 ft. how far did the car skid ?

OpenStudy (anonymous):

214 @phi

OpenStudy (phi):

yes, the car skidded 214 feet. Next, d is the distance that the car skids what number is d ?

OpenStudy (anonymous):

d=214

OpenStudy (phi):

yes. now we use that info. in \[ s = \sqrt{24d} \] replace d with 214 in that formula. (that means erase the d and put in 214) remember that 24d is short for 24 * d (24 times d)

OpenStudy (anonymous):



OpenStudy (phi):

almost. you start with \[ \sqrt{24d}\] replace the "d" with 214. take the "d" out (gone!) and put in 214 instead (like a substation of a player in a game)

OpenStudy (phi):

replace the "d" with 214. take the "d" out (gone!) and put in 214 instead (like a substitute player in a game)

OpenStudy (anonymous):

\[\sqrt{214} d24\]

OpenStudy (phi):

ok but take out the d

OpenStudy (anonymous):

\[\sqrt{214}^{24}\]

OpenStudy (phi):

yes, but it looks like this \[ \sqrt{24\cdot 214}\] the numbers are multiplied (in algebra we don't put in the multiplication sign, but that is what we mean) now multiply the numbers, then take the square root can you do that ?

OpenStudy (anonymous):

71.6658914687 @phi

OpenStudy (phi):

yes. the last step Round your answer to the nearest tenth.

OpenStudy (anonymous):

my answer choices are A. 6503.6 B. 7523.7 C. 2402.5 D. 2231.4 so the answe rwould be B

OpenStudy (phi):

those must be to a different question. A car does not go seven thousand miles per hour.

OpenStudy (anonymous):

OHH YES YES IT WAS I FOUND IT THANK YOU

OpenStudy (phi):

you got 71.6658914687 Round your answer to the nearest tenth. the "tenths" place is the 6 next to the decimal point to round, look at the number to its right (the other 6) if 5 or bigger (and it is), round the tenths place up 1 (to 7) the final answer is 71.7 rounded to the nearest tenth